# Position: AI agents should be regulated based on autonomous action sequences

### v1

@@ 4.3 Action trees, action graphs, and beyond
While an action sequence can well represent the behavior of a single AI agent, it may not be suitable when multiple AI agents are involved. Suppose, for example, that an AI agent takes a sequence of several actions and gives the results of taking those actions to another AI agent, and then these two agents sequentially take further actions in parallel to jointly accomplish a goal. In this scenario, the behavior of each agent can be represented by an action sequence, but these action sequences should not be considered independently. Rather, we should represent the behavior of the two agents with a tree of actions. In more complex scenarios, we would need to consider a graph of actions taken by multiple agents.
The similarity between action graphs may be defined analogously to action sequences. For example, we may say that any action graph of size $N$ is acceptable if any action graph of size $N-1$ is strongly acceptable, empirically.
@@
When a central controller coordinates multiple AI agents, it can track and manage the behaviors of those agents, making sure that the graph of their actions remain acceptable. When multiple AI agents collaborate in a decentralized manner, they should communicate their actions to each other to ensure that their action graph remain acceptable. In both cases, an action graph consists of actions taken autonomously by those agents. A difficulty arises when multiple AI agents collaborate implicitly by chance, and we currently do not have a solution to this.
4.4 On similarity-based measures
All the measures considered are based on similarity: we say that anything is acceptable when it is sufficiently close to something that is strongly acceptable. In Appendix B.1, we provide some rationale for these similarity-based measures.
When we say that a solution is acceptable based on a similarity measure, it is based on the assumptions that can be summarized with a tuple $({\mathcal{Z}}_{0},d,\varepsilon)$. Namely, what solutions are assumed to be strongly acceptable (${\mathcal{Z}}_{0}$), how the similarity between solutions is measured ($d$), and what level of guarantee is made ($\varepsilon$). For example, ${\mathcal{Z}}_{0}$ may be set of all the FMs that are trained with at most $10^{24}$ FLOPs of computation, $d$ may be the difference in FLOPs measured in $\log_{10}$, and $\varepsilon=1$. Alternatively, ${\mathcal{Z}}$ may be the set of action sequences of length at most 1000, $d$ may be the difference in the length of the action-sequences, and $\varepsilon=1$.
Governments may then design regulations based on the tuple $({\mathcal{Z}}_{0},d,\varepsilon)$. Scientists may provide guidelines regarding what tuples should be used in regulations.
@@ 5 Discussion and conclusion
We have argued that advanced AI agents should be regulated based on their action sequences or more generally action graphs. More broadly, we highlight the insufficiency of regulations that focus on the computational resources used for training AI models. Given the trend of increasing emphasis on sophisticated reasoning at inference time, we contend that the capabilities and existential risks of AI agents will stem largely from their inference-time computation.
@@
A key motivation for action sequences is the fundamental challenge of defining, representing, and observing the state of the open-ended environments in which AI agents operate. This limitation renders existing impact measures inadequate for managing existential risks. Since we cannot exhaustively identify all potential direct causes of global catastrophes or human extinction, it is impossible to determine a set of environmental features sufficient to predict whether a given action poses an existential risk in a specific state.
@@
The simplest form of the regulation based on action sequences would impose limits on their length, prohibiting AI agents from autonomously operating beyond the limit where their safety against existential risks is known empirically. This is neither a novel idea nor a sufficient solution; it should be seen as a baseline and a call to action for further research on this critical issue. Action sequences can allow more elaborate regulations, but more research is needed to better understand what they can or cannot guarantee. Further research is also needed to address myriad of questions, including what to be allowed as a single action, how to manage complex scenarios where multiple AI agents are involved, and how to ensure that the agents stop acting when their action sequence reaches the limit.
@@
The suggested approach of empirically verifying the strong acceptability of action sequences would inevitably delay the expansion of the range in which AI agents can operate autonomously. This is our intention. Although AI agents might acquire general or even superhuman intelligence with all the common sense that humans have, they can still pose existential risks, similar to humans. Humans, however, can often correct mistakes before they escalate, since our limited speed and scale allow us to identify and address errors—though there have been rare exceptions, such as those that led to world wars. In contrast, AI agents can operate at far greater speed and scale, preventing them from learning from mistakes in a controlled manner. Therefore, regulations should be seen not as obstacles to innovation and technological progress but as guidelines that accelerate research on making AI agents more controllable, verifiable, and governable, ensuring they truly benefit our society and the future.
Impact Statement
This paper proposes a new framework on regulating advanced AI agents, who can pose existential risks. The paper is thus intended to mitigate the negative impacts and implications that AI technologies might have. However, the paper also discusses how those advanced AI agents might be developed. While all the discussion is based on existing technologies that are already known in public, the unified treatment in this paper might motivate malicious readers to develop and deploy such advanced AI agents without sufficient considerations of their safety.
Appendix A Similarity between AI agents
It would be better if we could say that an AI agent is acceptable when it is sufficiently close to an AI agent that we empirically know is strongly acceptable, regardless of how those agents are developed (i.e., independent of their configurations including the amount of computation for training). A question is how to evaluate the similarity between two AI agents. If the two AI agents consist of identical base FMs and identical planning modules, their difference may be simply characterized by the amount of computation for inference. However, AI agents can consist of different base FMs or different planning modules.
@@ Appendix A Similarity between AI agents
A possible approach is to measure the similarity between two AI agents, $G$ and $G^{\prime}$, based on their output. For example, using a distance $d$ between probability distributions, we may for example define the distance between $G$ and $G^{\prime}$ with $\sup_{x}d(g(G,x),g(G^{\prime},x))$, where $g(G,x)$ is the distribution of the output (e.g., document) of $G$ when it perceives $x$ (e.g., prompt). This is similar to the motivation of reinforcement learning from human feedback Ouyang et al. (2022); Bai et al. (2022) and direct policy optimization Rafailov et al. (2023), where the regularization with Kullback–Leibler divergence is used to mitigate catastrophic forgetting or alignment tax Kotha et al. (2024); Ouyang et al. (2022); Bai et al. (2022), which refer to the phenomena that fine-tuned models lose the skills that pre-trained models had.
@@
Likewise, we may mitigate catastrophic forgetting of the strong acceptability of an AI agent $G^{\prime}$ and retain the acceptability in a new AI agent $G$ by ensuring $\sup_{x}d(g(G,x),g(G^{\prime},x))\leq\varepsilon$. A difficulty is that it is unclear how to evaluate the supremum, since there can be infinitely many possible perceptions $x$. Alternatively, one may consider ${\mathbb{E}}[d(g(G,X),g(G^{\prime},X))]$, where ${\mathbb{E}}$ is the expectation with respect to some distribution of the random perception $X$. However, such a guarantee based on expectation may be unsuitable for addressing safety concerns related to existential risks, which involve events with extremely low probabilities and extremely high impacts.
Appendix B Expanding the strongly acceptable set
This expansion of acceptable set is similar in spirit to safe exploration in RL García et al. (2015). Here, the safety set is gradually expand, starting from a seed set, for example based on the assumption of Lipshitz continuity and Gaussian process Sui et al. (2015). In control theory, safety is often guaranteed with barrier certificates often based on some prior knowledge about the environment Ames et al. (2019); Luo & Ma (2021); Bansal et al. (2017). Such ideas have also been exploited in safe exploration in RL with generative modeling Wang et al. (2023).
B.1 Rationale for similarity-based measures
@@
Here, we provide some rationale on such similarity-based measures. Let a solution $z$ denote a configuration, an AI agent, or an action-sequence; we discuss the acceptability of $z$.
@@
Appendix C Unobservable Markov decision processes
@@
There has been a limited amount of work on unobservable Markov decision processes (UMDPs) [^7] UMDPs have also been studied under the name of Markov decision processes (MDPs) with no observations, non-observable MDPs (NOMDPs), and no observation MDPs (NOMDPs).. An UMDP is a special case of a partially observable Markov decision process (POMDP) in that the agent always makes a null observation in the UMDP. A standard approach to finding the optimal policy for a POMDP is to recursively compute its value as a function of the belief state, which is updated on the basis of the Bayes rule. This is substantially simplified when there are no observations. In this section, we provide a comprehensive survey of the prior work on UMDPs.
@@
Madani et al. (1999; 2003) establish the undecidability of some decision problems associated with POMDPs over infinite horizons by establishing undecidability for the special case of UMDPs. Such undecidability for UMDPs can be established by reducing an UMDP to a probabilistic finite-state automaton. The undecidability also holds for a restricted class of UMDPs Balle et al. (2017; 2022). While approximate decision problems are still undecidable for general UMDPs over infinite horizons, Chatterjee et al. (2024) study a special case of UMDPs whose approximate decision problems are decidable.
@@
Burago et al. (1996) prove that computing the optimal policy for a POMDP over a finite horizon is NP-hard but showing that it is NP-hard for an UMDP. Wu & He (2020) study special cases of UMDPs over finite horizons whose optimal policies can be computed in polynomial time.
@@
UMDPs have also been used as approximations of POMDPs Hauskrecht (2000); Brechtel (2015); Lauri (2016) or simply discussed as a special case of POMDPs V (2009). For a given POMDP, the corresponding UMDP can given a lower bound on the value function, while the corresponding (fully observable) MDP can give an upper bound on the value function. This relation between UMDP and POMDP can be exploited to efficiently find approximately optimal policies for POMDPs. Notice that an UMDP can allow more efficient optimization than the corresponding POMDP, since the UMDP does not need to deal with observations. For exemple, King (2018) studies an MCTS method for UMDPs.
@@
UMDPs have also been studied as a simple special case of POMDPs to study the effectiveness of planning methods in belief states to study the relative performance of different planning methods for POMDPs Littlefield (2020); Littlefield et al. (2018); Kimmel et al. (2019). UMDPs have also been simply discussed as a special case of POMDPs Boutilier et al. (1999); Csáji (2008); Verma & Rao (2005).
@@
UMDP also appears in the study of planning for multiple distributed agents to optimize a single objective under partial observability. Specifically, planning for a decentralized POMDP (DecPOMDP) Oliehoek & Amato (2016) can be reduced to planning for a (centralized) UMDP Oliehoek & Amato (2014); Roijers (2016); Roijers & Whiteson (2020), where the state in the UMDP is the pair of the state and the history of observations in the DecPOMDP, and the action in the UMDP is the decision rule that maps the history of observations into the actions in the POMDP. In DecPOMDPs, the belief (distribution) over the pair of the state and the history of observations is the sufficient statistic, and planning can be performed in the space of the belief states.
@@
Even-Dar et al. (2007) consider an UMDP in the context of studying what values observations can provide in a POMDP. They introduce a parameter that ranges from 0 to 1. When the parameter is 0, the observation provides no information about the state (hence, the POMDP reduces to an UMDP); when the parameter is 1, the observation provides full information about the state (hence, the POMDP reduces to an MDP). The prior work also extends UMDPs to allow often costly actions that enable partial or full observations of the state Fox & Tennenholtz (2007); Kamar et al. (2009); Kamar (2010); Wang et al. (2025).

### v2

@@ 4.3 Action trees, action graphs, and beyond
While an action sequence may effectively represent the behavior of a single agent, it becomes insufficient in multi-agent scenarios. For example, one agent might perform a series of actions and pass the results to another agent, after which both agents may continue acting—possibly in parallel—to collaboratively achieve a shared goal. In such cases, although each agent’s behavior can be described by an individual action sequence, these sequences should not be analyzed in isolation. Instead, the joint behavior should be represented as an action tree that captures the interdependencies between agents. In more complex settings, this structure naturally generalizes to an action graph.
To assess the acceptability of such multi-agent behaviors, we can extend the notion of similarity from action sequences to action graphs. For instance, we might define an action graph of size $N$ as acceptable if all action graphs of size $N-1$ are strongly acceptable. Again, the size may be determined by counting only those actions in the designated subspace.
@@
When a central controller coordinates multiple agents, it can track and manage the behaviors of those agents, making sure that the graph of their actions remain acceptable. When multiple agents collaborate in a decentralized manner, they should communicate their actions to each other to ensure that their action graph remain acceptable. A difficulty arises when multiple agents collaborate implicitly by chance, and we currently do not have a solution to this.
4.4 Implementation and enforcement
While there are multiple ways to implement the proposed framework, one possible regulatory approach would require AI developers or deployers to submit the following information.
i)   Action Space: A description of the set of actions the AI agent is permitted to take autonomously.
ii)  Autonomy Limit: The maximum allowable length of autonomous action sequences before human
     intervention is required, along with a clear explanation of how this length is defined and measured. iii) Safety Evidence: Empirical evidence supporting the safety of shorter action sequences, including
     historical performance data, validation tests, and any relevant safety evaluations.
A regulatory authority could then review these submissions and make deployment decisions based on a structured risk assessment process.
Since developers and deployers define both the action space and the metric for action sequence length, they are naturally incentivized to define them in ways that enable agents to perform useful tasks. This design process effectively requires explicitly or implicitly enumerating all permissible actions—an essential foundation for ensuring safety. If an agent’s action space is so open-ended that it becomes intractable to ensure safety of individual actions in the action space, such an agent is not yet fit for deployment. Even when individual actions designed to be deemed safe, certain sequences of actions may still lead to existential risks. Our proposal specifically targets these risks by focusing on the regulation of autonomous action sequences.
We envision a framework in which the permissible action space and autonomy limit are defined for each type of AI agent. Some agents may require access to high-impact APIs—such as those used to control physical devices—while others may primarily focus on generating multi-modal contents. These agents differ significantly in their capabilities and potential risks, and therefore should be assigned distinct action spaces and autonomy limits accordingly.
For enforcement, the AI agent itself could be designed to track its own actions and terminate once it reaches the predefined cap. This could be achieved by integrating the action-sequence limit into the agent’s objective function, ensuring that it either completes its goal within the allowed length of action-sequence or shuts down automatically, similar to shut-down seeking agents proposed in [49].
Obviously, this approach cannot prevent malicious entities from exploiting technical loopholes to circumvent regulations. However, enforcement can still be achieved from a legal perspective, holding developers and deployers accountable for noncompliance and imposing penalties for violations.
@@ 5 Discussion and conclusion
We have argued that advanced AI agents should be regulated based on their action sequences or more generally action graphs. More broadly, we highlight the insufficiency of regulatory approaches that focus solely on training compute—particularly given the current trend of increasing emphasis on sophisticated reasoning at inference time—and advocate for complementary frameworks based on limiting the extent of autonomous operations.
@@
A central motivation for focusing on action sequences is the fundamental challenge of defining, representing, and observing the state of the open-ended environments in which agents operate. This limitation undermines the effectiveness of existing impact measures in managing existential risks. Since we cannot exhaustively identify all potential direct causes of global catastrophes or human extinction, it is infeasible to specify a set of environmental features that would reliably indicate whether a given action poses an existential risk in a particular state.
@@
The simplest form of the regulation based on action sequences would impose limits on their length, prohibiting AI agents from autonomously operating beyond the limit where their safety against existential risks is known empirically. This should be seen as a baseline and a call to action for further research on this critical issue. While action sequences offer a foundation for more nuanced regulatory frameworks, their full potential and limitations remain to be explored.
Key open questions include: What actions should be included in the designated subspace for measuring the length of action sequences? How should we manage complex scenarios involving multiple interacting agents? How can we ensure that agents reliably halt when reaching their autonomy limit? Addressing these challenges will be essential for developing robust safety standards.
@@
The proposed approach—empirically verifying the strong acceptability of action sequences—will inevitably slow the expansion of autonomous capabilities in AI agents. This is our intention. Even if AI agents attain general or superhuman intelligence with human-like common sense, they can pose existential risks, similar to humans. While humans can often correct mistakes before they escalate, thanks to our limited speed and scale, rare exceptions have led to catastrophic outcomes, such as world wars. In contrast, AI agents can operate at far greater speed and scale, preventing them from learning from mistakes in a controlled manner. Regulations should thus be seen not as obstacles to innovation and technological progress but as guidelines that accelerate research on making AI agents more controllable, verifiable, and governable, ensuring they truly benefit our society and the future.
Appendix A Biden’s executive order and related bills
In October 2023, Joe Biden, then president of the US, signed the Executive Order on the Safe, Secure, and Trustworthy Development and Use of Artificial Intelligence [18] [^5] This executive order was repealed by President Trump.. Its Section 4.2 is dedicated to ensuring safe and reliable AI. In particular, it requires companies to report on “any model that was trained using a quantity of computing power greater than $10^{26}$ integer or [FLOPs]” until a set of technical conditions for models are defined by specified authorities (see Section 4.2(b)).
In this executive order, particular attention is paid to a dual-use FM, which refers to an FM that exhibits “high levels of performance at tasks that pose a serious risk to security, national economic security, national public health or safety, or any combination of those matters, such as by … permitting the evasion of human control or oversight through means of deception or obfuscation” (see Section 3(k)).
Following this executive order, almost 700 AI-related bills are introduced in 45 states across the United States in 2024 [121]. A particularly interesting one is California Senate Bill 1047 (Safe and Secure Innovation for Frontier Artificial Intelligence Models Act) [136] [^6] The bill had passed the state legislature but was later vetoed by the Governor. The technical feasibility of the requirements has also been questioned by the community [2].. Its Chapter 22.6 is devoted to safe and secure innovation for frontier AI models, which cover “[a]n artificial intelligence model trained using a quantity of computing power greater than $10^{26}$ integer or floating-point operations.” In particular, the senate bill requires that “[b]efore beginning to initially train a covered model, the developer shall … [i]mplement the capability to promptly enact a full shutdown,” which completely halts the operations of the model.
The necessity of such an off-switch [53] is motivated to prevent “critical harms,” which include “[m]ass casualties or at least five hundred million dollars (\$500,000,000) of damage resulting from an artificial intelligence model engaging in conduct that … [a]cts with limited human oversight, intervention, or supervision.”
Appendix B Similarity between AI agents
In this section, we explore a method for determining whether an AI agent is acceptable, independent of how the agent is constructed—including factors such as training compute or architectural configuration—without relying on a detailed examination of its individual outputs. We may say that an agent is acceptable when it is sufficiently similar to another agent that has been empirically validated as strongly acceptable. A question is how to evaluate the similarity between two agents. If the two agents consist of identical base FMs and identical planning modules, their differences may be attributed to variations in inference compute (e.g., the number of search steps performed in MCTS). However, in practice, AI agents can vary widely in their base models, reasoning strategies, and system architectures. This diversity means that a single threshold based on inference compute is unlikely to serve as a universal criterion for acceptability. Also, it is unclear what constitute a single run of inference, as we have discussed in Section 3.2.
@@ Appendix B Similarity between AI agents
A possible approach is to measure the similarity between two agents, $G$ and $G^{\prime}$, based on their output distributions. For example, using a distance $d$ between probability distributions, we may for example define the distance between $G$ and $G^{\prime}$ with $\sup_{x}d(g(G,x),g(G^{\prime},x))$, where $g(G,x)$ is the distribution of the output (e.g., document) of $G$ when it perceives $x$ (e.g., prompt). This is similar to the motivation of reinforcement learning from human feedback [97, 11] and direct policy optimization [105], where the regularization with Kullback–Leibler divergence is used to mitigate catastrophic forgetting or alignment tax [75, 97, 11], which refer to the phenomena that fine-tuned models lose the skills that pre-trained models had.
@@
Likewise, we may mitigate catastrophic forgetting of the strong acceptability of an agent $G^{\prime}$ and preserve the acceptability in a new agent $G$ by ensuring $\sup_{x}d(g(G,x),g(G^{\prime},x))\leq\varepsilon$. A difficulty is that it is unclear how to evaluate the supremum, since there can be infinitely many possible perceptions $x$. Alternatively, one may consider ${\mathbb{E}}[d(g(G,X),g(G^{\prime},X))]$, where ${\mathbb{E}}$ is the expectation with respect to some distribution of the random perception $X$. However, such a guarantee based on expectation may be unsuitable for addressing safety concerns related to existential risks, which involve events with extremely low probabilities and extremely high impacts.
Appendix C Rationale for similarity-based measures
@@
All the measures that are considered in this paper are more or less based on similarity: we say that anything is acceptable when it is sufficiently close to something that is strongly acceptable. Here, we provide some rationale on such similarity-based measures. Let a solution $z$ denote a configuration, an AI agent, or an action-sequence; we discuss the acceptability of $z$.
@@
When we say that a solution is acceptable based on a similarity measure, it is based on the assumptions that can be summarized with a tuple $({\mathcal{Z}}_{0},d,\varepsilon)$. Namely, what solutions are assumed to be strongly acceptable (${\mathcal{Z}}_{0}$), how the similarity between solutions is measured ($d$), and what level of guarantee is made ($\varepsilon$). For example, ${\mathcal{Z}}_{0}$ may be set of all the FMs that are trained with at most $10^{24}$ FLOPs of computation, $d$ may be the difference in FLOPs measured in $\log_{10}$, and $\varepsilon=1$. Alternatively, ${\mathcal{Z}}$ may be the set of action sequences of length at most 1000, $d$ may be the difference in the length of the action-sequences, and $\varepsilon=1$. Governments may then design regulations based on the tuple $({\mathcal{Z}}_{0},d,\varepsilon)$. Scientists may provide guidelines regarding what tuples should be used in regulations.
Mathematical guarantees, such as safety under Lipschitz continuity, provide useful design principles for real-world systems. However, real-world safety cannot always be directly derived from such mathematical models, as assumptions in these models may not fully capture the complexity of actual systems. Even in highly regulated domains like aviation and nuclear safety, failures occasionally occur despite adherence to rigorous safety guidelines.
In our approach, we rely on the following key assumptions:
  • Any single action does not lead to catastrophic failure (e.g., human extinction) as long as it is taken from a strongly acceptable state.
  • An action-sequence of length $N-1$ is strongly acceptable and leads to a strongly acceptable state.
Here, a strongly acceptable state refers to a state that is reachable via strongly acceptable action sequences, starting from any initial states that normally occur in the absence of the agent. Based on these assumptions, we reason about the safety of an action sequence of length $N$ by ensuring that the $N$-th action is chosen from a strongly acceptable state. The validity of the assumptions depends on verification through empirical studies rather than theoretical guarantees.
We acknowledge that these assumptions do not hold for arbitrary action spaces. For instance, if the action space includes the action of launching a nuclear weapon, then a single action could lead to catastrophe. As we have discussed in Section 4.2, AI agents should be designed without such high-risk actions in their action space, and regulatory frameworks should prohibit AI agents that have access to actions with individually catastrophic consequences. More broadly, the design of the action space must be carefully structured to ensure practical safety.
Appendix D Expanding the strongly acceptable set
This expansion of acceptable set is similar in spirit to safe exploration in RL [47]. Here, the safety set is gradually expand, starting from a seed set, for example based on the assumption of Lipshitz continuity and Gaussian process [117]. In control theory, safety is often guaranteed with barrier certificates often based on some prior knowledge about the environment [4, 86, 14]. Such ideas have also been exploited in safe exploration in RL with generative modeling [132].
Appendix E Unobservable Markov decision processes
@@
The proposed approach may be considered as a way to ensure safety of actions in a Markov decision process where states are nearly unobservable (see Figure 2). In this section, we provide a comprehensive survey on a related model of unobservable Markov decision processes (UMDPs) [^8] UMDPs have also been studied under the name of Markov decision processes (MDPs) with no observations, non-observable MDPs (NOMDPs), and no observation MDPs (NOMDPs)., for which there has been a limited amount of prior work.
Figure 2: Safety of actions in a nearly unobservable Markov decision process
An UMDP is a special case of a partially observable Markov decision process (POMDP) in that the agent always makes a null observation in the UMDP. A standard approach to finding the optimal policy for a POMDP is to recursively compute its value as a function of the belief state, which is updated on the basis of the Bayes rule. This is substantially simplified when there are no observations. In this section, we provide a comprehensive survey of the prior work on UMDPs.
@@
[87, 88] establish the undecidability of some decision problems associated with POMDPs over infinite horizons by establishing undecidability for the special case of UMDPs. Such undecidability for UMDPs can be established by reducing an UMDP to a probabilistic finite-state automaton. The undecidability also holds for a restricted class of UMDPs [12, 13]. While approximate decision problems are still undecidable for general UMDPs over infinite horizons, [29] study a special case of UMDPs whose approximate decision problems are decidable.
@@
[26] prove that computing the optimal policy for a POMDP over a finite horizon is NP-hard but showing that it is NP-hard for an UMDP. [139] study special cases of UMDPs over finite horizons whose optimal policies can be computed in polynomial time.
@@
UMDPs have also been used as approximations of POMDPs [54, 24, 79] or simply discussed as a special case of POMDPs [125]. For a given POMDP, the corresponding UMDP can given a lower bound on the value function, while the corresponding (fully observable) MDP can give an upper bound on the value function. This relation between UMDP and POMDP can be exploited to efficiently find approximately optimal policies for POMDPs. Notice that an UMDP can allow more efficient optimization than the corresponding POMDP, since the UMDP does not need to deal with observations. For exemple, [73] studies an MCTS method for UMDPs.
@@
UMDPs have also been studied as a simple special case of POMDPs to study the effectiveness of planning methods in belief states to study the relative performance of different planning methods for POMDPs [81, 82, 72]. UMDPs have also been simply discussed as a special case of POMDPs [22, 34, 126].
@@
UMDP also appears in the study of planning for multiple distributed agents to optimize a single objective under partial observability. Specifically, planning for a decentralized POMDP (DecPOMDP) [93] can be reduced to planning for a (centralized) UMDP [92, 108, 109], where the state in the UMDP is the pair of the state and the history of observations in the DecPOMDP, and the action in the UMDP is the decision rule that maps the history of observations into the actions in the POMDP. In DecPOMDPs, the belief (distribution) over the pair of the state and the history of observations is the sufficient statistic, and planning can be performed in the space of the belief states.
@@
[41] consider an UMDP in the context of studying what values observations can provide in a POMDP. They introduce a parameter that ranges from 0 to 1. When the parameter is 0, the observation provides no information about the state (hence, the POMDP reduces to an UMDP); when the parameter is 1, the observation provides full information about the state (hence, the POMDP reduces to an MDP). The prior work also extends UMDPs to allow often costly actions that enable partial or full observations of the state [43, 68, 69, 130].